# The continuous spectrum of bound states in expulsive potentials

### v1

@@ Abstract
On the contrary to the common intuition that a steep expulsive potential makes quantum states widely delocalized, we demonstrate that one- and two-dimensional (1D and 2D) Schrödinger equations, which include expulsive potentials that are steeper than the quadratic (anti-harmonic-oscillator) ones, give rise to normalizable (effectively localized) eigenstates. These states constitute full continuous spectra in the 1D and 2D cases alike. In 1D, these are spatially even and odd eigenstates. The 2D states may carry any value of the vorticity (alias magnetic quantum number). Asymptotic approximations for wave functions of the 1D and 2D eigenstates, valid far from the center, are derived analytically, demonstrating excellent agreement with numerically found counterparts. Special exact solutions for vortex states are obtained in the 2D case. These findings suggest an extension of the concept of bound states in the continuum, in quantum mechanics and paraxial photonics. Gross-Pitaevskii equations are considered as the nonlinear extension of the 1D and 2D settings. In 1D, the cubic nonlinearity slightly deforms the eigenstates, maintaining their stability. On the other hand, the quintic self-focusing term, which occurs in the photonic version of the 1D model, initiates the dynamical collapse of states whose norm exceeds a critical value.
I Introduction
The separation of discrete and continuous spectra, which represent bound and delocalized states, respectively, is an underlying principle of the classical Sturm - Liouville theory [1, 2] and quantum mechanics, whose mathematical framework is based in this theory [3, 4, 5]. The fact that, in the paraxial approximation, the fundamental propagation equation for optical waves is tantamount to the quantum-mechanical Schrödinger equation makes the separation of discrete and continuous spectra an equally important tenet of optics and photonics [6, 7, 8]. Actually, this principle applies to various other physical settings modeled by equations of the Schrödinger type [9, 10].
Nevertheless, there are well-known exceptions from the separation principle, in the form of bound states embedded in the continuum (often designated by the BIC acronym). The first example was discovered in 1929 by von Neumann and Wigner (vNW) [11], solving the three-dimensional (3D) Schrödinger equation with the isotropic expulsive potential (written in a scaled form),
@@ I Introduction
$\varphi_{\mathrm{vNW}}(r)=\frac{1}{r^{2}}\sin\left(r^{3}\right),$ (2)
which corresponds to the zero energy eigenvalue. Later, schemes were designed for the realization of BIC states in semiconductor heterostructures [12, 13, 14]. Then, much interest was drawn to BIC modes in various photonic setups, where they offer a considerable potential for applications [15, 16, 17, 18, 19].
In the framework of the mean-field approximation, the linear Schrödinger equation for a single atom is closely related to its nonlinear version in the form of the Gross-Pitaevskii equation (GPE), which includes the cubic term accounting for the averaged effect of inter-atomic collisions in the Bose-Einstein-condensate (BEC) phase of ultracold atomic gases [21]. In the optical realization, a model similar to GPE is provided by the nonlinear Schrödinger equation (NLSE), which adds the Kerr self-focusing term to the paraxial propagation equation [6]. Under special conditions, NLSE-based optics models give rise to nonlinear states in the form of embedded solitons, which, similar to the BIC modes produced by the linear Schrödinger equation, may exist as peculiar localized states embedded in the continuous spectrum [22].
Another counter-intuitive species of bound states is produced by a system of two linearly coupled 1D or 2D equations, each one being either a linear or nonlinear Schrödinger equation, with the usual trapping harmonic-oscillator (HO) potential in one equation, and the expulsive (anti-HO) potential in the other [23]. The 2D linear system for wave functions $u$ and $v$ , written in the polar coordinates $\left(r,\theta\right)$, is
@@
with integer vorticity $S=0,1,2,$..., real eigenvalue $E$ (alias the chemical potential, in terms of BEC), which is a function of the system’s parameters $\lambda$, $\omega$, $\kappa$ and vorticity $S$ (as a special example, see Eq. (8) below), and stationary real wavefunctions are localized as usual eigenstates of the 2D HO, i.e., $\left\{U(r),V(r)\right\}\sim r^{S}\exp\left(-r^{2}/2\right)$ at $r\rightarrow\infty$, despite the presence of the expulsive potential in Eq. (4). In particular, if the system’s parameters are subject to the constraint $\omega=(1/2)\left(5+S-\lambda^{2}\right)$, the bound state is found as an exact solution,
@@
$E_{\mathrm{exact}}=\frac{1}{2}\left(\lambda^{2}+1+S\right)$ (8)
and with $U_{0}$ being an arbitrary amplitude. These strongly localized states, which exist with the single eigenvalue $E$, also represent BIC, as the same system gives rise to the continuous spectrum of weakly delocalized eigenstates, which are similar to the one given by Eq. (25) below.
The subject of the present work is to construct eigenstates of 1D and 2D linear and nonlinear Schrödinger equations (alias GPEs) with expulsive potentials, taken as the anti-HO term ($\gamma=1$ in Eq. (9)) or steeper ones ($\gamma>1$; note that the vNW example (1) includes $\gamma=2$). The respective 1D GPE for wave function $\psi\left(x,t\right)$ is introduced, in the scaled form, as
@@
where the coefficient in front of the potential term is fixed by means of scaling (cf. Eq. (3)), g=1,0,+1g=-1,0,+1 correspond to the self-focusing, zero, or defocusing nonlinearity, and two relevant values, $\sigma=2$ and $4$, represent, respectively, the cubic and quintic self-interaction. The cubic nonlinearity is commonly known as the most generic one [6], while the quintic self-focusing, which can be implemented [24, 25] and used [26, 27, 28, 29] under well-controlled conditions in optics [26, 24, 25], is interesting as it gives rise to Townes solitons in 1D [30]. Eigenstates produced by Eq. (9) with real eigenvalue $E$ are looked for in the usual form, $\psi\left(x,t\right)=\exp\left(-iEt\right)\varphi(x)$, with stationary real eigenfunction $\varphi(x)$ satisfying the equation
@@
In the experiments with cold atoms, the expulsive anti-HO or steeper potential, in its 1D and 2D forms alike, can be readily imposed by a blue-detuned optical beam with the Gaussian or super-Gaussian transverse structure [32, 31]. Another possibility of the experimental realization of the 1D model represented by Eq. (9) may refer to a gas of small polar molecules carrying a permanent electric dipole moment, cr. Refs. [33, 34]. Although this realization may seem too complex, one can consider the molecular gas loaded into a quasi-1D pipe-shaped optical trap, aligned with the $x$-axis, the dipoles being polarized in the positive and negative directions at $x>0$ and $x<0$, respectively, by a strong uniform electric field generated by a uniformly charged sheet placed at $x=0$ perpendicular to the $x$-axis. The expulsive potential $\sim 1/x^{2}$ ($\gamma=1$ in Eq. (9)) is then imposed by a point-like charge additionally placed at $x=0$, with the sign opposite to that that of the polarizing sheet.
@@
One may intuitively expect that the strongly expulsive potential in Eq. (9) leads to strong delocalization of the respective eigenstates. Nevertheless, we demonstrate, analytically and numerically, that the intuition is misleading: the 1D and 2D linear Schrödinger equations with the anti-HO potential, $\gamma=1$, give rise to weakly delocalized states (whose norm diverges logarithmically with the increase of the spatial size), while all steeper expulsive potentials, corresponding to $\gamma>1$, produces well-localized bound states with a convergent norm, which form a continuous spectrum ($\infty<E<+\infty$), This situation seems unusual in quantum mechanics.
While the most essential results are reported here in the framework of the 1D and 2D linear Schrödinger equations, numerical analysis demonstrates that the 1D bound states remain completely stable in the presence of the cubic term in Eq. (9) ($\sigma=1$). The quintic self-focusing term ($\sigma=2$ in Eq. (9), with g=1g=-1) gives rise to the collapse of the 1D bound states if their norm exceeds a certain critical value, while they are stable below the critical level. This situation is typical for the Townes-soliton phenomenology [35, 36, 37, 38].
@@
Analytical and numerical results for the bound states in the 1D linear Schrödinger equation are reported in Section 2. This is followed by a brief consideration of nonlinear bound states and their stability in the framework of the 1D NLSEs with the cubic and quintic self-focusing terms (g=1g=-1 and $\sigma=1$ or $2$, in terms of Eq. (9)), in Section 3. Findings for the 2D bound states, including vortex ones, which carry the angular momentum, are presented in Section 4 (chiefly, for the linear Schrödinger equation). In particular, the consideration of the 2D linear Schrödinger equation with the expulsive potential produces exact solutions for the vortex states, under a special constraint imposed on the model’s parameters. The paper is concluded by Section 5.
II One-dimensional linear bound states
@@
The starting point of the analysis is the linear version ($g=0$) of the 1D stationary equation (10). While the equation is not an analytically solvable one, it is straightforward to construct its asymptotic solution, valid for sufficiently large values of $|x|$ and $\gamma>1$:
$\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x;\gamma>1;E)=\varphi_{0}|x|^{-\gamma/2}\cos\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right)+\frac{E\varphi_{0}}{\gamma-1}|x|^{-3\gamma/2+1}\sin\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right),$ (11)
@@
where $\varphi_{0}$ and $\chi_{0}$ are arbitrary constants, in terms of the asymptotic expansion (note that setting $\gamma=1/2$ in the first term in Eq. (11) reproduces the well-known asymptotic form of the Airy function [3]). The approximation (11) retains the first two terms of the expansion, while its third-order correction ($\mathrm{TOC}$) reads
$\mathrm{TOC}\left(\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x);\gamma>1\right)=\frac{1}{8}\frac{\gamma\left(\gamma+2\right)}{\gamma+1}\varphi_{0}|x|^{-3\gamma/2-1}\sin\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right).$ (12)
@@
The obvious property of Eq. (11) is the convergence of its integral norm, $N=\int_{-\infty}^{+\infty}\varphi^{2}(x)dx$ at $|x|\rightarrow\infty$, for $\gamma>1$ and all eigenvalues (the full spectrum), $\infty<E<+\infty$. Thus, we arrive at a simple but counter-intuitive conclusion: any expulsive potential which is steeper than the anti-HO one ($\gamma>1$) produces normalizable bound states, which populate the full continuous spectrum. Qualitatively, this conclusion can be explained by noting that a classical counterpart of the quantum particle would be rolling down the steep potential hill with rapidly growing acceleration, which gives rise to the strongly oscillating phase of the respective wave function, the latter feature leading to the effective localization (normalizability) of the eigenstate.
@@
The situation is different in the case of the anti-HO potential, with $\gamma=1$ in Eq. (9). In this case, the approximation (11) is valid solely for $E=0$, yielding $\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x;\gamma=1;E=0)=\varphi_{0}|x|^{-1/2}\cos\left(x^{2}/2-\chi_{0}\right)$, demonstrating that the respective normalization integral slowly diverges,
$N\simeq\varphi_{0}^{2}\ln\left(L/l\right),$ (13)
where $2L$ is the size of the integration domain, and $l$ is a characteristic scale of the inner core of the wave function. For $E\neq 0$, the correct asymptotic approximation takes the form of
$\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x;\gamma=1;E)=\varphi_{0}|x|^{-1/2}\cos\left(\frac{x^{2}}{2}+E\ln\left(\frac{|x|}{l}\right)\right),$ (14)
whose norm diverges as in Eq. (13).
The wave function of the 1D eigenstate, produced by a numerical solution of Eq. (10) with $g=0$, $\gamma=1$ and $E=0$, and its comparison to the asymptotic approximation (11), are plotted in Fig. 1, with fitting constants $\chi_{0}=\pi/8$ and $\varphi_{0}=1$ in Eq. (11) (the latter value is actually arbitrarily for the linear solution). It is clearly observed that the asymptotic approximation provides high accuracy, in comparison to its numerical counterpart.
Figure 1: The continuous curve: a numerically found spatially even solution of Eq. (10) with $g=0$, $\gamma=1$ (the anti-HO expulsive potential) and $E=0$. The dashed curve: the asymptotic approximation (11) for the same solution, with fitting constants $\chi_{0}=\pi/8$ and $\varphi_{0}=1$.
@@
In Fig. 1, and similar in the numerically generated plots presented below, it is not easy to display far tails, as the rapidly accelerating phase oscillations in Eq. (11) imply that the accurate numerical solution must take into regard spatial harmonics of very high orders. On the other hand, the asymptotic expression (11) yields the tails in the virtually exact analytical form.
A typical example of the 1D bound state for $\gamma=2$ (the quartic expulsive potential), produced by the numerical solution of Eq. (10), and its comparison to the asymptotic approximation (11), is displayed in Fig. 2. In this case too, the asymptotic approximation demonstrates remarkably high accuracy.
@@
Figure 2: The continuous curve: a numerically found spatially even solution of Eq. (10) with $g=0$, $\gamma=2$ (the quartic expulsive potential) and $E=0$. The dashed curve: the asymptotic approximation (11) for the same case, with fitting constants $\chi_{0}=\pi/6$ and $\varphi_{0}=0.44$.
Note: For visualization purposes, the wavefunctions in several numerical figures were vertically rescaled. Accordingly, the plotted amplitudes do not represent the original normalization of the stationary solutions, and any amplitude-related fitting parameters (such as $\varphi_{0}$) correspond to the rescaled profiles shown in the plots.
For given $E$, the asymptotic-tail expressions (11) and (12) may belong, at least, to two different global solutions (full eigenstates), viz., the fundamental spatially even one, with $\varphi(-x)=\varphi(x)$ (alias the ground state, for given $E$), and the dipole spatially odd eigenstate, with $\varphi(-x)=-\varphi(x)$ (alias the first excited state). To illustrate this possibility for the same cases as considered above for the fundamental solutions, viz., $\gamma=1$ and $\gamma=2$, the corresponding numerically found dipole eigenstates, obtained with the boundary condition $d\varphi/dx|_{x=0}=0$, are plotted in Figs. 3(a) and (b), respectively. It is plausible that Eq. (10) may produce additional spatially even and odd solutions, which correspond to higher-order excited states in terms of quantum mechanics. This issue will be considered elsewhere.
@@
Figure 3: In panels (a) and (b), the continuous curves represent numerically found spatially odd (dipole-mode) solutions of Eq. (10) with $g=0$ and $E=0$, for $\gamma=1$ and $2$, respectively (cf. their spatially even counterparts in Figs. 1 and 2). The dashed curves represent the corresponding asymptotic approximation (11) with fitting constants $\chi_{0}=3\pi/8$, $\varphi_{0}=1.45$ in (a), and $\chi_{0}=\pi/3$, $\varphi_{0}=1.72$ in (b).
@@
While the above examples display eigenstates found with the zero eigenvalue ($E=0$), Fig. 4 presents a typical eigenstate for a very large negative eigenvalue, viz., E=200E=-200, in the case of $\gamma=2$ (the quartic expulsive potential). It is seen that the large value of E-E suppresses the wave function in the core area of the eigenstate. As the energy decreases, the wave function goes away from the potential maximum around $x=0$. As $E\rightarrow$-$\infty$, $x_{\max}\rightarrow\infty$ and the wave function approaches the ground state. At the edge of the core, the wave function quickly grows towards the largest value, attained at a point with $|x|=x_{\max}$, as $\varphi(x)\sim\exp\left(\sqrt{-2E}\left(x-x_{\max}\right)\right)$, in agreement with Eq. (10). This is followed by the gradual decay of the rapidly oscillating tail, as per Eq. (11). The dependence of $x_{\max}$ on $|E|$ may be predicted, in a crude approximation, by equating the two respective large terms in Eq. (10), i.e., $E\varphi\sim(1/2)x_{\max}^{4}\varphi$, hence, in the logarithmic approximation,
@@
Figure 5: The numerically found dependence of the coordinate $x_{\max}$, at which the wave function attains the largest value, on the eigenvalue $E<0$, for $g=0$ and $\gamma=2$. The dependence is plotted on the log-log scale, with the straight dashed line corresponding to the approximate relation (15).
@@
In the eigenstates corresponding to very large positive eigenvalues, the core area, $|x|\lesssim E^{1/4}$, is not (nearly) empty, unlike Fig. 5. Instead, it is filled by a uniform standing wave, $\varphi(x)\approx\varphi(x=0)\cos\left(\sqrt{2E}x\right)$ (not shown here in detail).
@@
III Bound states in the 1D nonlinear Schrödinger equation (NLSE) with the expulsive potential

### v2

@@ Abstract
On the contrary to the common intuition, which suggests that a steep expulsive potential makes quantum states widely delocalized, we demonstrate that one- and two-dimensional (1D and 2D) Schrödinger equations, which include expulsive potentials that are steeper than the quadratic ones, give rise to normalizable eigenstates, which may be considered as a manifestation of effective self-trapping in the linear system. These states constitute full continuous spectra in both the 1D and 2D cases. In 1D, they are spatially even and odd eigenstates. The 2D states may carry any value of the vorticity (alias magnetic quantum number). Asymptotic expressions for wave functions of the 1D and 2D eigenstates, valid far from the center, are derived analytically, demonstrating excellent agreement with full numerical solutions. Special exact solutions for vortex states are obtained in the 2D case. These findings suggest an extension of the concept of bound states in the continuum, in quantum mechanics and paraxial photonics. Gross-Pitaevskii equations are briefly considered as the nonlinear extension of the 1D and 2D settings. In 1D, the cubic nonlinearity slightly deforms the eigenstates, maintaining their stability
Keywords: bound states in continuum; self-trapping; normalizability; asymptotic structure; cubic nonlinearity; parity-breaking instability; vorticity; angular momentum
I Introduction and methods
The separation of discrete and continuous spectra, which represent bound and delocalized states, respectively, is an underlying principle of the classical Sturm - Liouville theory [1, 2, 3, 4] and quantum mechanics, whose mathematical framework is based on this theory [5, 6]. The fact that, in the paraxial approximation, the fundamental propagation equation for optical waves is tantamount to the quantum-mechanical Schrödinger equation makes the separation of discrete and continuous spectra an equally important tenet of optics [7, 8, 9, 10, 11]. This principle applies as well to other physical settings modeled by equations of the Schrödinger type [12, 13].
Nevertheless, there are well-known exceptions from the separation principle, in the form of bound states embedded in the continuum (often designated by the BIC acronym). The first example was discovered in 1929 by von Neumann and Wigner (vNW) [14], as a solution of the three-dimensional (3D) Schrödinger equation with the isotropic expulsive potential (written here in a scaled form),
@@ I Introduction and methods
$\varphi_{\mathrm{vNW}}(r)=\frac{1}{r^{2}}\sin\left(r^{3}\right)$ (2)
(see Fig. 1), which corresponds to the zero energy eigenvalue. Later, schemes were designed for the realization of quantum-mechanical BIC states in semiconductor heterostructures [15, 16]. Then, much interest was drawn to BIC modes in various photonic setups, where they offer a considerable potential for applications [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Figure 1: The wave function $\varphi_{\mathrm{vNW}}(r)$ of the von Neumann-Wigner BIC state.
In the framework of the mean-field approximation, the linear Schrödinger equation for a single atom is closely related to its nonlinear version in the form of the Gross-Pitaevskii equation (GPE), which includes the cubic term accounting for the averaged effect of inter-atomic collisions in the Bose-Einstein-condensate (BEC) phase of ultracold atomic gases [28]. In the optical realization, a model similar to GPE is based on the nonlinear Schrödinger equation (NLSE), which adds the Kerr self-focusing term to the paraxial propagation equation [7]. Under special conditions, NLSE-based optical models give rise to nonlinear states in the form of embedded solitons, which, similar to the BIC modes produced by the linear Schrödinger equation, may exist as peculiar localized states embedded in the continuous spectrum [29, 30, 31, 32, 33, 34].
Another counter-intuitive species of bound states is produced by a system of two linearly coupled 1D or 2D equations, each one being either a linear or nonlinear Schrödinger equation, with the usual trapping harmonic-oscillator (HO) potential in one equation, and the expulsive (anti-HO; alias inverted HO [35]) potential in the other [36]. The 2D linear system for wave functions $u$ and $v$ , written in the polar coordinates $\left(r,\theta\right)$, is
@@
with integer vorticity $S=0,1,2,$..., real eigenvalue $E$ (alias the chemical potential, in terms of BEC), which is a function of the system’s parameters $\lambda$, $\omega$, $\kappa$ and vorticity $S$ (as a special example, see Eq. (8) below), and stationary real wavefunctions which are localized as usual eigenstates of the 2D HO, i.e., $\left\{U(r),V(r)\right\}\sim r^{S}\exp\left(-r^{2}/2\right)$ at $r\rightarrow\infty$, despite the presence of the expulsive potential in Eq. (4). In particular, if the system’s parameters are subject to constraint $\omega=(1/2)\left(5+S-\lambda^{2}\right)$, the bound state is found as an exact solution,
@@
$E_{\mathrm{exact}}=\frac{1}{2}\left(\lambda^{2}+1+S\right),$ (8)
where $U_{0}$ is an arbitrary amplitude. These strongly localized states, which exist with the single eigenvalue $E$, also represent BIC, as the same system gives rise to the continuous spectrum of weakly delocalized eigenstates, which are similar to the one given by Eq. (31) below.
The subject of the present work is to construct eigenstates of 1D and 2D linear and (in a brief form) nonlinear Schrödinger equations (alias GPEs) with expulsive potentials, taken as the quadratic anti-HO term ($\gamma=1$ in Eq. (9)) or steeper ones ($\gamma>1$; note that the vNW example (1) includes $\gamma=2$). The respective 1D GPE for wave function $\psi\left(x,t\right)$ is introduced, in the scaled form, as
@@
where the coefficient in front of the potential term is fixed by means of scaling (cf. Eq. (3)), g=1,0,+1g=-1,0,+1 correspond to the self-focusing, zero, or defocusing nonlinearity, and two physically relevant values, $\sigma=1$ and $2$, represent, respectively, the cubic and quintic self-interaction. The cubic nonlinearity is the most generic one [7], while the quintic self-focusing, which can be implemented [37] and used [38] under well-controlled conditions in optics [38, 37], is interesting, as it gives rise to the 1D variety [39] of Townes solitons [40, 41, 41]. Eigenstates produced by Eq. (9) with real eigenvalue $E$ are looked for in the usual form, $\psi\left(x,t\right)=\exp\left(-iEt\right)\varphi(x)$, with stationary real eigenfunction $\varphi(x)$ satisfying the equation
@@
In the experiments with cold atoms, the expulsive anti-HO (quadratic) or steeper potential, in its 1D and 2D forms alike, can be readily imposed by a blue-detuned optical beam with the properly shaped transverse structure [43, 44, 45], or a red-tuned one with intrinsic vorticity [46, 47, 48]. The recently developed technique of programmable optical potentials [49, 50, 51, 52] can also be used for this purpose.
@@
One may intuitively expect that the strongly expulsive potential in Eq. (9) leads to strong delocalization of the respective eigenstates. Nevertheless, we demonstrate, analytically and numerically, that the intuition is misleading: the 1D and 2D linear Schrödinger equations with the anti-HO (quadratic) potential, $\gamma=1$, give rise to weakly delocalized states, whose norm,
$N_{\mathrm{1D}}=\int_{-\infty}^{+\infty}\varphi^{2}(x)dx,$ (11)
diverges logarithmically with the increase of the size of the solution domain (see Eq. (15) below), while all steeper expulsive potentials, corresponding to $\gamma>1$, produce effectively localized (self-trapped) bound states with a convergent norm, which form a continuous spectrum ($\infty<E<+\infty$), This situation seems unusual in quantum mechanics and BEC theory (as sort of a quantum anomaly [54, 35, 55, 53, 56]) – in particular, because self-trapping is usually considered as the feature which may only be induced by the nonlinearity, while here it takes place in the linear setting.
While the most essential results are reported here in the framework of the 1D and 2D linear Schrödinger equations, some results for the 1D NLSE are reported, in a brief form, too. It is demonstrated that numerically found 1D bound states remain stable or develop parity-breaking instability, below or above a critical value of the norm $N_{\mathrm{1D}}$, respectively, under the action of the cubic self-focusing term ($\sigma=1$, g=1g=-1) in Eq. (9) with the quartic expulsive potential ($\gamma=2$).
@@
Analytical and numerical results for the bound states in the 1D linear Schrödinger equation are reported in subsection 2A. This is followed, in subsection 2B, by a brief consideration of nonlinear bound states and their stability in the above-mentioned case, with parameters $\sigma=1$, g=1g=-1, and $\gamma=2$ in Eq. (9),Findings for the 2D bound states, including vortex ones, which carry the angular momentum, are presented in subsection 2C (chiefly, for the linear Schrödinger equation). In particular, the consideration of the 2D linear Schrödinger equation with the expulsive potential produces exact solutions for the vortex states, under a special constraint imposed on the model’s parameters. The paper is concluded by Section 3.
II Results
II.1 One-dimensional linear bound states
@@
The starting point of the analysis is the linear version ($g=0$) of the 1D stationary equation (10). While the equation is not analytically solvable, it is straightforward to construct its asymptotic solution, valid for sufficiently large values of $|x|$, in the form of a combination of rapidly oscillating cosines and sines, with amplitudes expanded in appropriate powers of $|x|^{-1}$. The analytical calculation is similar to the commonly known one which produces the asymptotic approximation for the Airy function [5]. Thus, the asymptotic approximation yields the result which is valid for $\gamma\neq 1$:
$\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x;\gamma\neq 1;E)=\varphi_{0}|x|^{-\gamma/2}\cos\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right)+\frac{E\varphi_{0}}{\gamma-1}|x|^{-3\gamma/2+1}\sin\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right),$ (12)
@@
where the amplitude $\varphi_{0}$ and phase shift $\chi_{0}$ are arbitrary constants, in terms of the asymptotic solution. Note that setting $\gamma=1/2$ in the first term in Eq. (12) reproduces the above-mentioned asymptotic approximation for the Airy function. The approximation (12) retains the first two terms of the expansion, the third-order correction ($\mathrm{TOC}$) being
$\mathrm{TOC}\left(\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x);\gamma\neq 1\right)=\frac{1}{8}\frac{\gamma\left(\gamma+2\right)}{\gamma+1}\varphi_{0}|x|^{-3\gamma/2-1}\sin\left(\frac{|x|^{\gamma+1}}{\gamma+1}-\chi_{0}\right).$ (13)
@@
An obvious property of the wave function (12) is the convergence of its integral norm, $N=\int_{-\infty}^{+\infty}\varphi^{2}(x)dx$ at $|x|\rightarrow\infty$, for $\gamma>1$ and all eigenvalues $\infty<E<+\infty$ (the full continuous spectrum). Thus, we arrive at a simple but counter-intuitive conclusion: any expulsive potential which is steeper than the quadratic (alias anti-HO) one, with $\gamma>1$, produces normalizable bound states, which populate the full continuous spectrum. Qualitatively, this conclusion can be explained by noting that a classical counterpart of the quantum particle would be rolling down the steep potential hill with the rapidly growing acceleration, which gives rise to a strongly oscillating phase in the respective wave function, the latter feature leading to the effective self-trapping (localization/normalizability) of the eigenstate in the linear system, unlike the commonly known mechanism of self-trapping (creation of solitons) in various nonlinear models [7].
@@
The situation is different in the case of the anti-HO potential, with $\gamma=1$ in Eq. (9). In this case, approximation (12) is replaced by the following one, which includes a logarithmic correction to the phase of the oscillatory wave function:
$\varphi_{\mathrm{asympt}}^{\mathrm{(1D)}}(x;\gamma=1)=\varphi_{0}|x|^{-1/2}\cos\left(\frac{x^{2}}{2}+E\ln\left(\frac{|x|}{l}\right)\right),$ (14)
where $l$ is a characteristic scale of the inner core of the wave function (an exact wave function of the 1D anti-HO potential can be formally expressed in terms of the special parabolic-cylinder function [35]). It follows from expression (14) that the respective normalization integral (11) slowly diverges,
$N_{\mathrm{1D}}\simeq\varphi_{0}^{2}\ln\left(L/l\right),$ (15)
$2L$ being the size of the integration domain.
The wave function of the 1D linear eigenstate, produced by a numerical solution of Eq. (10) with $g=0$, $\gamma=1$ (the anti-HO potential) and $E=0$, and its comparison to the asymptotic approximation (12), are plotted in Fig. 2, with fitting constants $\chi_{0}=\pi/8$ and $\varphi_{0}=1$ in Eq. (12) (in fact, $\varphi_{0}$ is an arbitrarily constant for the linear solution). The respective numerical solution of the second-order ordinary differential equation (10) was produced by means of the Runge-Kutta method, with the initial conditions
$\varphi(x=0)=1,\frac{d\varphi}{dx}(x=0)=0$ (16)
(which means that even solutions, with $\varphi(-x)=\varphi(x)$, were sought for). The same numerical method was used to obtain various eigenstates in what follows below. It is clearly observed that the analytical asymptotic approximation provides high accuracy, in comparison to its numerical counterpart.
@@
In Fig. 2 it is not easy to display slowly decaying tails of the wave function, as the rapidly accelerating phase oscillations in Eq. (12) imply that the accurate numerical solution must take into regard spatial harmonics of very high orders. On the other hand, the asymptotic expression (12) yields the tails in the virtually exact analytical form.
@@
Figure 2: The continuous curve: the numerically found spatially even solution of Eq. (10) with $g=0$, $\gamma=1$ (the quadratic expulsive potential) and $E=0$. The dashed curve: the asymptotic approximation (12) for the same solution, with fitting constants $\chi_{0}=\pi/8$ and $\varphi_{0}=1$.
A typical example of the normalizable (effectively self-trapped) 1D bound state for $\gamma=2$ (the quartic expulsive potential), produced by the numerical solution of Eq. (10), and its comparison to the asymptotic approximation (12), is displayed in Fig. 3. In this case too, the asymptotic approximation demonstrates remarkably high accuracy.
@@
Figure 3: The continuous curve: the numerically found spatially even solution of Eq. (10) with $g=0$, $\gamma=2$ (the quartic expulsive potential) and $E=0$. The dashed curve: the asymptotic approximation (12) for the same case, with fitting constants $\chi_{0}=\pi/6$ and $\varphi_{0}=0.44$.
For given $E$, the asymptotic expressions (12) and (13) may belong, at least, to two different global solutions (full eigenstates), viz., the fundamental spatially even one, with $\varphi(-x)=\varphi(x)$ (alias the ground state, for given $E$), and the dipole spatially odd eigenstate, with $\varphi(-x)=-\varphi(x)$ (alias the first excited state). To illustrate this possibility for the same cases as considered above for the fundamental solutions, viz., $\gamma=1$ and $\gamma=2$, the corresponding numerically found dipole eigenstates are plotted in Figs. 4(a) and (b), respectively. To this end, Eq. (10) was numerically solved again by means of the Runge-Kutta method, with the initial conditions
$\varphi(x=0)=0,~\frac{d\varphi}{dx}\left(x=0\right)=1,$ (17)
cf. Eq. (16). It is plausible that Eq. (10) may produce additional spatially even and odd solutions, which correspond to higher-order excited states, in terms of quantum mechanics. This possibility will be considered elsewhere.
Figure 4: In panels (a) and (b), the continuous curves represent numerically found spatially odd (dipole-mode) solutions of Eq. (10) with $g=0$ and $E=0$, for $\gamma=1$ and $2$, respectively (cf. their spatially even (fundamental) counterparts displayed in Figs. 2 and 3, resectively). The dashed curves represent the corresponding asymptotic approximation (12) with fitting constants $\chi_{0}=3\pi/8$, $\varphi_{0}=1.45$ in (a), and $\chi_{0}=\pi/3$, $\varphi_{0}=1.72$ in (b).
@@
While the above examples display eigenstates found with the zero eigenvalue ($E=0$), Fig. 5 presents a typical eigenstate for a very large negative eigenvalue, viz., E=200E=-200, in the case of $\gamma=2$ (the quartic expulsive potential). It is seen that the large value of E-E suppresses the wave function in the core area of the eigenstate. At the inner edge of the core, the wave function quickly grows towards its largest value, attained at a point with $|x|=x_{\max}$, as $\varphi(x)\sim\exp\left(\sqrt{-2E}\left(x-x_{\max}\right)\right)$, in agreement with Eq. (10). This is followed by the gradual decay of the rapidly oscillating tail, in agreement with Eq. (12). The dependence of $x_{\max}$ on $|E|$ may be predicted, in a crude approximation, by equating the two respective large terms in Eq. (10), i.e., $E\varphi\sim(1/2)x_{\max}^{4}\varphi$ (recall we here set $\gamma=2$), which yields, in the logarithmic approximation,
@@
Figure 6: The numerically found dependence of the coordinate $x_{\max}$, at which the wave function attains its maximum, on the eigenvalue $E<0$, for $g=0$ and $\gamma=2$. The dependence is plotted on the log-log scale, with the straight dashed line representing the approximate relation (18).
@@
In the eigenstates corresponding to very large positive eigenvalues, the core area, $|x|\lesssim E^{1/4}$, is not (nearly) empty, unlike Fig. 6. Instead, it is filled by a nearly uniform standing wave, $\varphi(x)\approx\varphi(x=0)\cos\left(\sqrt{2E}x\right)$ (not shown here in detail).
@@
II.2 Stable and unstable bound states in the 1D cubic nonlinear Schrödinger equation (NLSE) with the expulsive potential